OpenStudy (anonymous):

Given f(x) = 4x + 1 all over 3 solve for f-1(3).

OpenStudy (anonymous):

2 4 6 8

OpenStudy (ybarrap):

$$ f(x)=\cfrac{4x+1}{3} $$ And you need, $$ f^{-1}(3) $$ That is you need the inverse evaluated at 3?

OpenStudy (anonymous):

im confused

OpenStudy (ybarrap):

Can you read the attached? Is this your problem?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

@ybarrap

OpenStudy (ybarrap):

ok, replace f(x) with x and x with y and solve for y.

OpenStudy (ybarrap):

That'll give you your inverse. Then plug in x=3 to the result. Done.

OpenStudy (anonymous):

let y=f(x)=(4x+1)/3 interchange x and y x=(4y+1)/3 4y=3x-1 y=1/4 (3x-1) f^-1(x)=1/4(3x-1) plug x=3 and get the solution.

OpenStudy (anonymous):

I got 2 @ybarrap and @surjithayer

OpenStudy (ybarrap):

That's right - (1/4) (3(3) -1 ) = (9-1)/4 =8/4 = 2